# Plentiful Jailbreaks with String Compositions

### v2

@@ 1.2 Related work
To reiterate, many encodings mentioned in the introduction, including leetspeak, Morse Code, low-resource language translations, rotary ciphers, and ASCII, fall under the purview of invertible transformations. Besides encodings, the adversarial attack literature for language models has included gradient-based discrete optimization (Zou et al., 2023; Liu et al., 2024; Shin et al., 2020; Ebrahimi et al., 2017; Guo et al., 2021; Geisler et al., 2024; Zhu et al., 2023; Guo et al., 2024; Thompson and Sklar, 2024); LLM-assisted prompt optimization (Chao et al., 2023; Mehrotra et al., 2023; Tang et al., 2024); multi-turn or many-shot attacks (bijection; Li et al., 2024; Russinovich et al., 2024; Anil et al., 2024; Zheng et al., 2024); and other idiosyncratic attack vectors (Andriushchenko and Flammarion, 2024; Andriushchenko et al., 2024).

### v3

@@ 1.2 Related work
To reiterate, many encodings mentioned in the introduction, including leetspeak, Morse Code, low-resource language translations, rotary ciphers, and ASCII, fall under the purview of invertible transformations. Besides encodings, the adversarial attack literature for language models has included gradient-based discrete optimization (Zou et al., 2023; Liu et al., 2024; Shin et al., 2020; Ebrahimi et al., 2017; Guo et al., 2021; Geisler et al., 2024; Zhu et al., 2023; Guo et al., 2024; Thompson and Sklar, 2024); LLM-assisted prompt optimization (Chao et al., 2023; Mehrotra et al., 2023; Tang et al., 2024); multi-turn or many-shot attacks (Li et al., 2024; Haize Labs, 2024; Russinovich et al., 2024; Anil et al., 2024; Zheng et al., 2024); and other idiosyncratic attack vectors (Huang et al., 2024; Andriushchenko and Flammarion, 2024; Andriushchenko et al., 2024).